# Are We Done with MMLU?

### v2

@@ Appendix B Prompting Methods
Standard Prompting # Task: Given a question, its choices, and the ground truth answer, classify the question as either ’ok’ or ’not ok’. - ’ok’ means that the question and the choices are understandable, and the ground truth answer is correct. - ’not ok’ means that the ground truth answer is incorrect, or the question and the choices are not well presented. Classify with ’ok’ or ’not ok’ WITHOUT PROVIDING ANY REASONING
Chain of Thought (CoT) Prompting # Task: Given a triple consisting of a multiple choice question, its choices, and the corresponding ground truth answer, your task is to classify the triple into ’ok’ or ’not ok’. # Instructions: 1. Question Presentation: Is the question well-presented? Assess clarity, grammar, and sufficiency of information. 1.1 If Yes, assess the presentation of the MC options. 1.2 If No, classify the issue as ’not ok’. 2. MC Options Presentation: Are the MC options well-presented? Check if the options are clear, distinct, and relevant to the question. 2.1 If Yes, determine if there is one potentially correct answer. 2.2 If No, classify the issue as ’not ok’. 3. Answer Evaluation: Is there one, more than one, or no potentially correct answer in the options list? 3.1 If one, continue to Ground Truth Answer Evaluation. 3.2 If more than one, classify the issue as ’not ok’. 3.3 If no correct answer, classify the issue as ’not ok’. 4. Ground Truth Answer Evaluation: Is the ground truth answer correct? 4.1. If Yes, classify as ok. 4.2. If No, classify as ’not ok’. Provide your assessment in JSON format with keys ’Question Presentation’, ’MC Options Presentation’, ’Answer Evaluation’, ’Ground Truth Answer Evaluation’, ’Classification’. The ’classification’ is either ok, or not ok. FOLLOW THE EXACT EXAMPLE ANSWER FORMAT ALL IN ONE LINE WITHOUT PROVIDING EXPLANATION # Example Answer: {"Question Presentation": "OK", "MC Options Presentation": "OK", "Answer Evaluation": "One", "Ground Truth Answer Evaluation": "Correct", "Classification": "OK"}
Appendix C Details on automatic error detection
C.1 Detailed Results on Error Detection Experiments for MMLU-Redux
Table 5: Comparison of Zero-Shot, Chain-of-Thought (CoT), and Few-Shot F1 score.
@@ C.1 Detailed Results on Error Detection Experiments for MMLU-Redux
Table 6: Comparison of Zero-Shot, Chain-of-Thought (CoT), and Few-Shot F2 score.
@@
Table 7: Comparison of Zero-Shot, Chain-of-Thought (CoT), and Few-Shot Recall score.
@@
Appendix D Heterogeneity of Errors in MMLU Subsets – Full List
Here, we provide an extensive list of qualitative observations from the manually validated MMLU subsets. Some of these observations were previously reported in by the AI Explained YouTube Channel [12] and a Medium blog post [13].
@@ Appendix D Heterogeneity of Errors in MMLU Subsets – Full List
– Incorrect ground truths labels are particularly prevalent within the Virology subset. Many of the incorrect labels are for relatively simple questions, such as identifying the description of a pandemic, this suggests errors may stem from problems parsing the original datasets (in most cases the Human Virology textbook’s student resources). In addition, there are a range of issues relating to question and option clarity where the necessary context required to answer the question is missing, such as which family of diseases is being referred to.
@@
– This subset includes several unclear questions where respondents must identify multiple correct statements in a listed statement. MMLU provides only the last statement of the list as the question instead of including the entire question.
@@
– This is almost completely correct.
@@
– This subset is almost completely correct. The majority of the questions are from AP Statistics and can be directly obtained from the ‘crackap’ website.
@@
– Some of the questions (approximately 20%) were duplicated.
@@
– This is almost completely correct. There are only a few questions (mostly basic equations) with wrong ground truth answer.
@@
Appendix E Error Detection via Fine-tuning
@@
To create LabelChaos, we selected and merged six manually labelled datasets. We chose datasets annotated by humans ([39, 40, 41, 42, 43]) to avoid the potential inaccuracies associated with automated labelling procedures. After standardising these datasets to align with the format of MMLU, we generate a corrupted version of each by introducing specific types of corruption, as categorised in our taxonomy (Fig. 3). These corruptions were carefully applied to replicate the quality and distribution characteristics of the MMLU dataset. The final dataset comprises approximately 264,000 samples, similar to those in MMLU. Below, we provide an overview of the LabelChaos subsets and the corruption procedures used for each type of error.
@@
We fine-tune the Llama-3 (8B-Instruct) [8] using LabelChaos datasets. To balance the distribution in MMLU-Redux, where most instances are labelled as "correct", we adjusted the label distribution to: 0.1 (Wrong Ground Truth), 0.1 (Poor Question Clarity), 0.1 (No Correct Answers), 0.1 (Unclear Options), 0.1 (Multiple Correct Answers), and 0.5 (correct). For consistency, we used the previously described CoT Prompt as input, with outputs labelled as either “ok” or “not ok”. The training involves 2048 steps, with a batch size of 64, utilizing the AdamW optimizer [44] with a learning rate of $2\times 10^{-4}$ and no weight decay. Due to computational constraints, we apply LoRA [45], with a rank of 16, to all models. All experiments were conducted on a single Nvidia A100 (40GB) GPU. We present the results of the fine-tuning method in Table 8.

### v3

@@ Appendix G Prompting Methods
Standard Prompting # Task: Given a question, its choices, and the ground truth answer, classify the question as either ‘ok’ or ‘not ok’. - ‘ok’ means that the question and the choices are understandable, and the ground truth answer is correct. - ‘not ok’ means that the ground truth answer is incorrect, or the question and the choices are not well presented. Classify with ‘ok’ or ‘not ok’ WITHOUT PROVIDING ANY REASONING
Chain of Thought (CoT) Prompting # Task: Given a triple consisting of a multiple choice question, its choices, and the corresponding ground truth answer, your task is to classify the triple into ‘ok’ or ‘not ok’. # Instructions: 1. Question Presentation: Is the question well-presented? Assess clarity, grammar, and sufficiency of information. 1.1 If Yes, assess the presentation of the MC options. 1.2 If No, classify the issue as ‘not ok’. 2. MC Options Presentation: Are the MC options well-presented? Check if the options are clear, distinct, and relevant to the question. 2.1 If Yes, determine if there is one potentially correct answer. 2.2 If No, classify the issue as ‘not ok’. 3. Answer Evaluation: Is there one, more than one, or no potentially correct answer in the options list? 3.1 If one, continue to Ground Truth Answer Evaluation. 3.2 If more than one, classify the issue as ‘not ok’. 3.3 If no correct answer, classify the issue as ‘not ok’. 4. Ground Truth Answer Evaluation: Is the ground truth answer correct? 4.1. If Yes, classify as ‘ok’. 4.2. If No, classify as ‘not ok’. Provide your assessment in JSON format with keys ‘Question Presentation’, ‘MC Options Presentation’, ‘Answer Evaluation’, ‘Ground Truth Answer Evaluation’, ‘Classification’. The ‘classification’ is either ‘ok’, or ‘not ok’. FOLLOW THE EXACT EXAMPLE ANSWER FORMAT ALL IN ONE LINE WITHOUT PROVIDING EXPLANATION # Example Answer: {“Question Presentation”: “Correct”, “MC Options Presentation”: “Correct”, “Answer Evaluation”: “One”, “Ground Truth Answer Evaluation”: “Correct”, “Classification”: “Correct”}
Appendix H Details on automatic error detection
H.1 Detailed Results on Error Detection Experiments for MMLU-Redux
Table 6: Comparison of Zero-Shot, Chain-of-Thought (CoT), and Few-Shot F1 score on 10 MMLU-Redux subjects.
@@ H.1 Detailed Results on Error Detection Experiments for MMLU-Redux
Table 7: Comparison of Zero-Shot, Chain-of-Thought (CoT), and Few-Shot F2 score on 10 MMLU-Redux subjects.
@@
Table 8: Comparison of Zero-Shot, Chain-of-Thought (CoT), and Few-Shot Recall score on 10 MMLU-Redux subjects.
@@
Appendix I Heterogeneity of Errors in MMLU Subsets – Full List
Here, we provide an extensive list of qualitative observations from the manually validated MMLU subsets. Some of these observations were previously reported in by the AI Explained YouTube Channel (L and Stapleton, 2023) and a Medium blog post (Erenrich, 2023).
@@ Appendix I Heterogeneity of Errors in MMLU Subsets – Full List
– Incorrect ground truths labels are particularly prevalent within the Virology subset. Many of the incorrect labels are for relatively simple questions, such as identifying the description of a pandemic, this suggests errors may stem from problems parsing the original datasets (in most cases the Human Virology textbook’s student resources). In addition, there is a range of issues relating to question and option clarity where the necessary context required to answer the question is missing, such as which family of diseases is being referred to.
@@
– This subset includes several unclear questions where respondents must identify multiple correct statements in a listed statement. MMLU provides only the last statement of the list as the question, instead of including the entire question.
@@
– This is almost entirely correct.
Abstract algebra
– This subset contains several answers that are simply incorrect. Some answers are ambiguous because they assume a ring does not need to have multiplicative identities, which is usually assumed in modern treatments of rings. This changes the outcome of the given answer.
@@
– This subset is almost entirely correct. The majority of the questions are from AP Statistics and can be directly obtained from the ‘crackap’ website.
@@
– Some questions (approximately 20%) were duplicated.
@@
– This is almost entirely correct. There are only a few questions (mostly basic equations) with wrong ground truth answers.
@@
Professional Psychology
– This subset mainly contains questions in the style of Examination for Professional Practice in Psychology (EPPP) mainly used in most U.S. states and Canadian provinces. Among 100 randomly sampled questions, there are four erroneous instances. For instance, one of them contains no correct answer due to incorrect parsing (i.e., incorrectly parsing ‘0.12’ as ‘0.42’). Another example is a bad question clarity where the answer is highly dependent on the geographical context (i.e., ‘In most cases, statutory responsibility for establishing minimal standards for professional competency to protect the public from harm rests with …’).
Security Studies
– The questions are sourced from textbooks (e.g.,Collins: Contemporary Security Studies 6e Student Resources’). Six out of 100 instances have bad question clarity due to underspecified abbreviations (e.g., ‘HM’ and ‘TNC’). When read in this form without other supporting context, readers may confuse the abbreviation with other relevant concepts in the security studies. In the textbook, the questions were presented as part of a specific chapter, and the abbreviations were defined in the earlier questions.
High School Computer Science
– This subset is completely clean.
High School Psychology
– The questions are mostly sourced from AP Psychology practice questions. Among the 100 instances, there are four problematic ones. Three of them are due to missing statements. These questions are presented as multiple true-false questions (i.e., the options are ‘I only’, ‘II only’, ‘III only’, and ‘I and II only’) without the four statements. One of the four erroneous instances is problematic due to racial stereotyping (i.e.,Aisha is a beautiful, black teenager. If she is typical, she most likely believes that’).
Moral Disputes
– This subset contains four erroneous instances. For instance, three of them refers to underspecified entities (e.g.,Ashford’s article is meant to address a particular paralysis in the face of’, ‘If Thomson’s conclusion is correct, then’, ‘In responding to terrorism, says Walzer, it is particularly important that’).
Moral Scenarios
– This subset contains three instances with ambiguous questions requiring more background information before deciding whether the statement is ‘morally wrong’.
High School World History
– The questions are sourced from publicly available AP World History practice questions from sources such as ‘crackap’. This subset is almost completely correct, except for one question with a wrong ground truth which may be caused by a random mistake.
High School European History
– The errors in this subset are mainly due to bad option clarity. In some cases, there is an option that is equivalent to the correct one. Additionally, in a few samples, the passage provided as a reference does not contain the answer to the question, despite the question implying that it should.
Computer Security
– The questions are sourced from university-level quizzes, with most of them being correct. There are three instances with wrong ground truth answer, however it is unclear what may have caused it.
Prehistory
– Completely clean.
US Foreign Policy
– The questions were mostly sourced from textbooks (e.g., Cox & Stokes: US Foreign Policy 3e) and publicly available final exams. There is only one question with bad clarity in terms of the temporal context (i.e., ‘How many states in the international system are likely to have nuclear weapons right now?’).
Sociology
– This subset contains questions that are sourced from textbooks (e.g.,Fulcher & Scott: Sociology 4e’). There are two instances with bad question clarity, mainly because the question requires time and geographical contexts (i.e.,Patterns of drug use in Britain reveal that:’ and ‘Which of the following policies did the New Labour government not pursue?’).
World Religions
– This subset is almost completely correct. One instance with a bad question clarity requires more context about when is this question asked (i.e.,When was the current Dalai Lama born?’).
College Biology
– This subset is mostly sourced from a publicly available GRE Biology practice test. There are two erroneous instances; one of them has a wrong ground truth answer, while the other one has its ground truth answer copy pasted to the question due to incorrect parsing (i.e.,Stabilization of the unique coiled structure of an alpha helix in a protein is primarily attributed to (A) hydrogen bonding between the peptide backbone atoms’).
High School Biology
– Most questions in this subset are sourced from AP Biology practice questions. There are five erroneous instances, three of them seem to be caused by bad parsing. For instance, the correct option A is simply written as ‘0’. Another erroneous example is where the question refers to a previous question (i.e.,Refer to question 15 for details on the squirrel population. Which of the following conditions is required to keep this population in Hardy-Weinberg equilibrium?’).
International Law
– Most questions in this subset are sourced from a textbook (i.e., ‘Bantekas & Papastavridis: International Law Concentrate 4e’), with most of them are correct.
Jurisprudence
– Most questions in this subset are sourced from a textbook (i.e., ‘Wacks: Understanding Jurisprudence 5e’), with most of them are correct.
Medical Genetics
– This subset is sourced from university-level final exams and is completely correct.
Professional Medicine
– This subset contains questions that are sourced from the publicly available sample questions of a standardised medical exam in the US (i.e., the United States Medical Licensing Examination (USMLE)). Several questions with an unknown source can be answered with several searches on established websites (e.g., National Center for Biotechnology Information (NCBI), Mayo Clinic). This subset only contains one erroneous instance due to bad question clarity where the original question has an accompanying photograph.
Marketing
– The questions in this subset are sourced from university-level quizzes and textbooks. There are four erroneous instances. One question misses the temporal context, while the other is resented as a statement with unclear relation with the options.
High School Microeconomics
– The questions in this subset are sourced from AP Microeconomics. Most of them are correct, except for two instances. One instance has a wrong ground truth answer for unknown reason. Another one has a bad question clarity, referring to unspecified question (i.e.,The basis for the answer in number 1 is’).
Management
– This is almost completely correct.
Nutrition
– This is almost entirely correct. Only one question has a minor mistake in one of its options (i.e., the ‘>’ sign was mistakenly replaced by ‘?’ mark) and one question was annotated by adding the time when the question was asked (similar to Clinical Knowledge) to indicate the ever-changing nature of the questions.
Human Sexuality
– Most questions in this subset are sourced from university-level exams that are publicly available. There are 17 erroneous instances of various types. Several questions are highly subjective or speculative (e.g.,Persons with liberal attitudes about premarital sex are likely to:’, ‘Alexander the Great was probably’ (in relation to his sexuality)).
High School Macroeconomics
– The questions in this subset are sourced from AP Macroeconomics. Most of them are correct, except for three instances. One instance has two identical correct answers. The other two are fill-in-the-blank type of questions, however the position of the blanks are not specified.
High School Government and Politics
– The questions are mostly sourced from AP U.S. Government and Politics practice questions, with most answers can be found on the ‘crackap’ website. This subset is completely correct.
High School US History
– The questions in this subset are sourced from AP U.S. History practice questions, with all answers can be found on the ‘crackap’ website. This subset is completely correct.
Appendix J Error Detection via Fine-tuning
@@
To create LabelChaos, we selected and merged six manually labelled datasets. We chose datasets annotated by humans ((Clark et al., 2018; Mihaylov et al., 2018; Amini et al., 2019; Jin et al., 2021; Lin et al., 2021b)) to avoid the potential inaccuracies associated with automated labelling procedures. After standardising these datasets to align with the format of MMLU, we generate a corrupted version of each by introducing specific types of corruption, as categorised in our annotation protocol (Fig. 3). These corruptions were carefully applied to replicate the quality and distribution characteristics of the MMLU dataset. The final dataset comprises approximately 264,000 samples, similar to those in MMLU. Below, we provide an overview of the LabelChaos subsets and the corruption procedures used for each type of error.
@@
We fine-tune the Llama-3 (8B-Instruct) (Dubey et al., 2024) using LabelChaos datasets. To balance the distribution in MMLU-Redux, where most instances are labelled as "correct", we adjusted the label distribution to: 0.1 (Wrong Ground Truth), 0.1 (Poor Question Clarity), 0.1 (No Correct Answers), 0.1 (Unclear Options), 0.1 (Multiple Correct Answers), and 0.5 (correct). For consistency, we used the previously described CoT Prompt as input, with outputs labelled as either “ok” or “not ok”. The training involves 2048 steps, with a batch size of 64, utilising the AdamW optimizer (Loshchilov and Hutter, 2018) with a learning rate of $2\times 10^{-4}$ and no weight decay. Due to computational constraints, we apply LoRA (Hu et al., 2021), with a rank of 16, to all models. All experiments were conducted on a single Nvidia A100 (40GB) GPU. We present the results of the fine-tuning method in Table 9.